You are given the following information for Smashville, Inc.

Cost of goods sold: $ 204,000
Investment income: $ 2,000
Net sales: $ 369,000
Operating expense: $ 80,000
Interest expense: $ 7,400
Dividends: $ 10,000
Tax rate: 40 %
Current liabilities: $ 16,000
Cash: $ 21,000
Long-term debt: $ 28,000
Other assets: $ 36,000
Fixed assets: $ 163,000
Other liabilities: $ 5,000
Investments: $ 40,000
Operating assets: $ 35,000

Calculate the gross margin, the operating margin, return on assets, and return on equity. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)
Answer #1
Net sales 369000
(-) Cost of goods sold 204000
Gross profit 165000
(-) Operating expense 80000
Operating income 85000
(+) Investment income 2000
(-) Interest expense 7400
Earnings before tax 79600
(-) Tax @ 40% 31840
Net income 47760
Fixed assets 163000
Cash 21000
Other assets 36000
Investment 40000
Operating assets 35000
Total assets 295000
Total assets 295000
(-) Current liabilities 16000
(-) Long term debt 28000
(-) Other liabilities 5000
Total equity 246000
Gross margin = Gross profit / Net sales = 165000 / 369000 44.72%
Operating margin = Operating income / Net sales = 85000 / 369000 23.04%
Return on assets = Net income / Total assets = 47760 / 295000 16.19%
Return on equity = Net income / Total equity = 47760 / 246000 19.41%